# Self-organized arrival system for urban air mobility

### v1

@@ Abstract
Urban air mobility is an innovative mode of transportation in which electric vertical takeoff and landing (eVTOL) vehicles operate between nodes called vertiports. We outline a self-organized vertiport arrival system based on deep reinforcement learning. The airspace around the vertiport is assumed to be circular, and the vehicles can freely operate inside. Each aircraft is considered an individual agent and follows a shared policy, resulting in decentralized actions that are based on local information. We investigate the development of the reinforcement learning policy during training and illustrate how the algorithm moves from suboptimal local holding patterns to a safe and efficient final policy. The latter is validated in simulation-based scenarios and also deployed on small-scale unmanned aerial vehicles to showcase its real-world usability.
@@ 1 Introduction
  • We conduct testing through scenario-based validation and a thorough simulation study. Moreover, we investigate and visualize how the policy changes during training, delivering further insights into the learning process.
  • Finally, we demonstrate the effectiveness of our policy, entirely trained in simulation, through real-world experiments conducted on small-scale unmanned aerial vehicles called Crazyflies (Giernacki et al., 2017). The demonstrated success of the Sim-2-Real transfer highlights the practical applicability of our results.
The paper is structured as follows: Section 2 reviews related work regarding safety assurance systems and terminal arrival organization in UAM operations. Section 3 defines the airspace design and describes the RL-based modeling approach. Subsequently, Section 4 details the definition of the observation, action, and reward spaces, the algorithmic background, and the training environment. Afterward, Section 5 displays the results, including the real-world experiments. Section 6 concludes this paper. The source code to this work is publicly available at the GitHub repository Waltz and Paulig (2022).
@@ 3.1 Airspace design
We consider a circle-shaped terminal airspace around a vertiport, which is visualized in Figure 1. The vertiport VTOL zone is assumed to have a radius of $200\,\mathrm{m}$, followed by an airspace with an outer radius of $R=800\,\mathrm{m}$. Following Bertram and Wei (2020b), the incoming traffic is assumed to enter the airspace through one of the four gates, which are located in the north, east, south, and west directions from the vertiport. Afterward, the aircraft can freely operate inside the airspace, resembling free-flight concepts of conventional aircraft traffic (Hoekstra et al., 2002; Ribeiro et al., 2022; Kleinbekman et al., 2018; Groot et al., 2024). The vehicles are unable to hover, possibly travel at different speeds, and can only adjust their respective heading. These assumptions, though restrictive, significantly elevate the complexity of the problem. However, contrary to Bertram and Wei (2020b), we do not impose further constraints like pre-defined traffic rings inside the airspace or particular traveling directions of individual vehicles. Hence, the clock-wise direction of the five shown aircraft shown in Figure 1 serves purely as an illustration. In addition, all vehicles operate on the same altitude to avoid interactions with departing aircraft at lower altitude.
@@ 4.2 Observation and action space
$o_{t}=\left(\left(o_{\text{O},t}\right)^{\top},\left(o_{\text{T},t}\right)^{\top}\right)^{\top},$
@@
$o_{\text{O},t}=\left(\frac{\left[\alpha_{\text{O},t}^{\text{V}}\right]^{\pi}_{-\pi}}{\pi},\frac{d_{\text{O},t}^{\text{V}}}{d_{\rm scale}},\sigma_{\text{O},t}\right)^{\top},$
@@
$\left[\theta\right]_{-\pi}^{\pi}=\begin{cases}\theta-\left\lfloor\frac{\theta+\pi}{2\pi}\right\rfloor\cdot 2\pi&\quad\text{if }\theta\geq 0,\\ \theta+\left\lfloor\frac{-\theta+\pi}{2\pi}\right\rfloor\cdot 2\pi&\quad\text{if }\theta<0,\\ \end{cases}$
@@
$o_{\text{T},i,t}=\left(\frac{d_{\text{O},t}^{\text{i}}}{d_{\rm scale}},\frac{\left[\alpha_{\text{O},t}^{\text{i}}\right]^{\pi}_{-\pi}}{\pi},\frac{v_{i,t}-v_{\text{O},t}}{v_{\rm scale}},\frac{\left[\psi_{i,t}-\psi_{\text{O},t}\right]^{\pi}_{-\pi}}{\pi},\frac{d_{t,\text{CPA}}^{i}}{\tilde{d}_{\rm scale}},\frac{t_{t,\text{CPA}}^{i}}{t_{\rm scale}}\right)^{\top},$
@@
$\psi_{\text{O},t+1}=[\psi_{\text{O},t}+a_{t}\cdot\Delta]_{-\pi}^{\pi},$
@@ 4.3 Reward function
$r_{\text{goal},t}=\begin{cases}(d_{\text{O},t-1}^{\text{V}}-d_{\text{O},t}^{\text{V}})/c_{4}&\text{for}\quad\sigma_{\text{O},t}=1,\\ -5&\text{for}\quad(\sigma_{\text{O},t}=-1)\land(d_{\text{O},t}^{\text{V}}\leq c_{5}),\\ 0&\rm else,\end{cases}$
@@
$r_{\text{space},t}=\begin{cases}-5&\text{for}\quad d_{\text{O},t}^{\text{V}}\geq 1000\,\mathrm{m},\\ 0&\rm else.\end{cases}$
@@
$r_{t}=\omega_{\text{coll}}\cdot r_{\text{coll},t}+\omega_{\text{goal}}\cdot r_{\text{goal},t}+\omega_{\text{space}}\cdot r_{\text{space},t}+\omega_{\text{comf}}\cdot r_{\text{comf},t},$
@@ 4.4 Algorithm
$z_{\mu,t-l}$ $=f_{\mu,l}\left(o_{\text{O},t-l},o_{\text{T},t-l};\theta_{f_{\mu,l}}\right)\quad\text{for}\quad l=0,\ldots,h,$
@@
$z_{j,t-l}$ $=f_{j,l}\left(o_{\text{O},t-l},o_{\text{T},t-l};\theta_{f_{j,l}}\right)\quad\text{for}\quad l=0,\ldots,h,$
@@ 5.2 Policy development
We sequentially initialize 30 vehicles, with the details being provided in Section 5.4. The results are presented in a publicly accessible video [^1] https://youtu.be/4475p8YqrEg.
@@ 5.4 Simulation study
We perform a thorough simulation study to quantitatively evaluate the Policy IV’s performance. In particular, we spawn in total ${N\in\{5,10,15,25,30\}}$ vehicles per episode, and the video provided in Section 5.1 illustrates the case for $N=30$. The aircraft appear at a random gate with a $15\,\mathrm{s}$ time gap. We disturb the heading of each aircraft, which points toward the eVTOL zone midpoint, by a realization of $\mathcal{U}(-20,20)\cdot 1^{\circ}$. In addition, we randomly sample the aircraft speeds from $\mathcal{U}(10,16)\cdot 1\,\mathrm{m/s}$. We repeat the experiment 30 times for each $N$ and track the number of accidents and the number of incidents as safety metrics. In addition, we measure the number of false vertiport entrances and track the average time of the aircraft needed to enter the goal after receiving the entrance signal, quantifying the efficiency of the operations.
@@
A more profound understanding of these events is gleaned from the video presented in Section 5.2. The root cause of most incidents lies in aircraft entering the space, particularly when other vehicles are near the gate. In an operational environment, this problem is solved by pre-arrival management, which informs aircraft already in the cruise phase about arrival times and provides advisories for in-time arrivals. Recognizing this observation, we conducted a secondary simulation study, incorporating a simple additional safety check for airspace entrance. Specifically, we examined whether any aircraft within the airspace maintained a Euclidean distance of less than $300\,\mathrm{m}$ to the respective gate. If affirmative, an alternative gate was selected for entrance; if all gates were occupied, no entrance is possible.
@@
In addition, Figures 7 and 8 show that there are almost no false vertiport entrances except in cases with higher traffic density. Moreover, the average entrance time is reduced from approximately $45\,\mathrm{s}$ on average for $N=5$ to slightly above $30\,\mathrm{s}$ for $N=30$ vehicles. This result is intuitive since a higher traffic density forces the vehicles to spread around the available airspace more broadly, thereby reducing the distance of the closest vehicle to the eVTOL zone.
@@ 5.5 Sim-2-Real transfer
We demonstrate the real-world applicability of our RL-based control approach by deploying Policy IV on five Crazyflie drones (Giernacki et al., 2017). The results are showcased in a publicly available video [^2] https://youtu.be/wbHUxARpHzM. As depicted in Figure 10, the Crazyflie is a nano quadcopter used as an experimental platform for research and education. In our experiment, we estimate the drones’ positions using a camera-based motion capture system, while the ROS2-based Crazyswarm2 (IMRCLab, 2024) testbed acts as middleware for communicating with the drones. Since the RL policy provides high-level heading commands, we compute the desired position of each drone assuming a constant velocity, as described in Section 3.1. This desired position serves as input for a differential flatness-based low-level controller (Mellinger and Kumar, 2011), which governs the drone actuators.
In our experimental setup, five drones are initially equidistantly spaced in the airspace. The radius of the airspace is set to $1.2\,\mathrm{m}$, and the drones maintain a velocity of approximately $0.15\,\mathrm{m/s}$. The video illustrates a stable and efficient operation, with the drones maintaining a safe distance and entering the vertiport sequentially. Crucially, we did not retrain the policy to accommodate the reduced size of the environment compared to the training scenarios. Since the agents’ observation space is properly normalized (see Section 4.2), adjusting the normalizing constants to match the smaller scale of the experiment suffices. The successful zero-shot adaptation of our policy is remarkable, considering that Sim-2-Real transfer often poses a significant challenge for RL algorithms (Zhao et al., 2020; Muratore et al., 2019).

### v2

# Self-organized free-flight arrival for urban air mobility

@@ Abstract
Urban air mobility is an innovative mode of transportation in which electric vertical takeoff and landing (eVTOL) vehicles operate between nodes called vertiports. We outline a self-organized vertiport arrival system based on deep reinforcement learning. The airspace around the vertiport is assumed to be circular, and the vehicles can freely operate inside. Each aircraft is considered an individual agent and follows a shared policy, resulting in decentralized actions that are based on local information. We investigate the development of the reinforcement learning policy during training and illustrate how the algorithm moves from suboptimal local holding patterns to a safe and efficient final policy. The latter is validated in simulation-based scenarios, including robustness analyses against sensor noise and a changing distribution of inbound traffic. Lastly, we deploy the final policy on small-scale unmanned aerial vehicles to showcase its real-world usability.
@@ 1 Introduction
  • We conduct testing through scenario-based validation, thorough simulation studies, and different robustness analyses. Moreover, we investigate and visualize how the policy changes during training, delivering further insights into the learning process.
  • Crucially, we demonstrate the effectiveness of our policy, entirely trained in simulation, through real-world experiments conducted on small-scale unmanned aerial vehicles called Crazyflies (Giernacki et al., 2017). The demonstrated success of the Sim-2-Real transfer highlights the practical applicability of our results.
The paper is structured as follows: Section 2 reviews related work regarding safety assurance systems and terminal arrival organization in UAM operations. Section 3 defines the airspace design and describes the RL-based modeling approach. Subsequently, Section 4 details the definition of the observation, action, and reward spaces, the algorithmic background, and the training environment. Afterward, Section 5 displays the results, while Section 6 contains the robustness analysis, including the real-world experiments. Section 7 concludes this paper. The source code to this work is publicly available at the GitHub repository Waltz and Paulig (2022).
@@ 3.1 Airspace design
We consider a circle-shaped terminal airspace around a vertiport, which is visualized in Figure 1. The vertiport VTOL zone is assumed to have a radius of $200\,\mathrm{m}$, followed by an airspace with an outer radius of $R=800\,\mathrm{m}$. As stated in European Union Aviation Safety Agency (2022), safety regions with obstacle clearance will surround vertiports. Hence, we consider the airspace to be free of static obstacles, which would impact the vehicles’ flight trajectories. Following Bertram and Wei (2020b), the incoming traffic is assumed to enter the airspace through one of the four gates, which are located in the north, east, south, and west directions from the vertiport. Afterward, the aircraft can freely operate inside the airspace, resembling free-flight concepts of conventional aircraft traffic (Hoekstra et al., 2002; Ribeiro et al., 2022; Kleinbekman et al., 2018; Groot et al., 2024). The vehicles are assumed to be unable to hover, possibly travel at different speeds, and can only adjust their respective heading. These assumptions, though restrictive, significantly elevate the complexity of the problem. In particular, while eVTOL vehicles typically have the ability to hover, the energy consumption is significantly higher than in the flight mode (Kasliwal et al., 2019), where horizontal speed additionally results in dynamic lift. From a conventional air traffic management perspective, our approach can thus be considered as a combination of holding patterns (International Civial Aviation Organization, 2018) and the point merge concept (semi-circle approach area; Eurocontrol 2021). In particular, our airspace design effectively adjusts the point merge concept to a circular area.
Contrary to Bertram and Wei (2020b), we do not impose further constraints like pre-defined traffic rings inside the airspace or particular traveling directions of individual vehicles. Hence, the clock-wise direction of the five shown aircraft shown in Figure 1 serves purely as an illustration. In addition, similar to conventional air traffic, we assume that flight levels will separate arrival and departure flows. In particular, all arrival vehicles operate on the same altitude, avoiding interactions with departing aircraft at lower altitude.
@@ 4.2 Observation and action space
$o_{t}=\left(\left(o_{\text{O},t}\right)^{\top},\left(o_{\text{T},t}\right)^{\top}\right)^{\top},$ (1)
@@
$o_{\text{O},t}=\left(\frac{\left[\alpha_{\text{O},t}^{\text{V}}\right]^{\pi}_{-\pi}}{\pi},\frac{d_{\text{O},t}^{\text{V}}}{d_{\rm scale}},\sigma_{\text{O},t}\right)^{\top},$ (2)
@@
$\left[\theta\right]_{-\pi}^{\pi}=\begin{cases}\theta-\left\lfloor\frac{\theta+\pi}{2\pi}\right\rfloor\cdot 2\pi&\quad\text{if }\theta\geq 0,\\ \theta+\left\lfloor\frac{-\theta+\pi}{2\pi}\right\rfloor\cdot 2\pi&\quad\text{if }\theta<0,\\ \end{cases}$ (3)
@@
$o_{\text{T},i,t}=\left(\frac{d_{\text{O},t}^{\text{i}}}{d_{\rm scale}},\frac{\left[\alpha_{\text{O},t}^{\text{i}}\right]^{\pi}_{-\pi}}{\pi},\frac{v_{i,t}-v_{\text{O},t}}{v_{\rm scale}},\frac{\left[\psi_{i,t}-\psi_{\text{O},t}\right]^{\pi}_{-\pi}}{\pi},\frac{d_{t,\text{CPA}}^{i}}{\tilde{d}_{\rm scale}},\frac{t_{t,\text{CPA}}^{i}}{t_{\rm scale}}\right)^{\top},$ (5)
@@
$\psi_{\text{O},t+1}=[\psi_{\text{O},t}+a_{t}\cdot\Delta]_{-\pi}^{\pi},$ (6)
@@ 4.3 Reward function
$r_{\text{goal},t}=\begin{cases}(d_{\text{O},t-1}^{\text{V}}-d_{\text{O},t}^{\text{V}})/c_{4}&\text{for}\quad\sigma_{\text{O},t}=1,\\ -5&\text{for}\quad(\sigma_{\text{O},t}=-1)\land(d_{\text{O},t}^{\text{V}}\leq c_{5}),\\ 0&\rm else,\end{cases}$ (8)
@@
$r_{\text{space},t}=\begin{cases}-5&\text{for}\quad d_{\text{O},t}^{\text{V}}\geq 1000\,\mathrm{m},\\ 0&\rm else.\end{cases}$ (9)
@@
$r_{t}=\omega_{\text{coll}}\cdot r_{\text{coll},t}+\omega_{\text{goal}}\cdot r_{\text{goal},t}+\omega_{\text{space}}\cdot r_{\text{space},t}+\omega_{\text{comf}}\cdot r_{\text{comf},t},$ (10)
@@ 4.4 Algorithm
$z_{\mu,t-l}$ $=f_{\mu,l}\left(o_{\text{O},t-l},o_{\text{T},t-l};\theta_{f_{\mu,l}}\right)\quad\text{for}\quad l=0,\ldots,h,$ (11)
@@
$z_{j,t-l}$ $=f_{j,l}\left(o_{\text{O},t-l},o_{\text{T},t-l};\theta_{f_{j,l}}\right)\quad\text{for}\quad l=0,\ldots,h,$ (12)
@@ 5.2 Policy development
We sequentially initialize 30 vehicles, with the details being provided in Section 5.4. The results are presented in a publicly accessible animation [^1] https://youtu.be/4475p8YqrEg.
@@ 5.4 Simulation study
We perform a thorough simulation study to quantitatively evaluate the Policy IV’s performance. In particular, we spawn in total ${N\in\{5,10,15,25,30\}}$ vehicles per episode, and the animation provided in Section 5.2 illustrates the case for $N=30$. The aircraft appear at a random gate with a $15\,\mathrm{s}$ time gap. We disturb the heading of each aircraft, which points toward the eVTOL zone midpoint, by a realization of $\mathcal{U}(-20,20)\cdot 1^{\circ}$. In addition, we randomly sample the aircraft speeds from $\mathcal{U}(10,16)\cdot 1\,\mathrm{m/s}$. We repeat the experiment 30 times for each $N$, tracking the number of accidents and incidents as safety metrics. Additionally, we monitor the number of false vertiport entrances. To quantify operational efficiency, we measure the average airspace time, defined as the duration from entering the airspace to entering the vertiport. This metric is the sum of two components: The average time to signal, which is the time it takes for an aircraft to receive the entrance signal, and the average entrance time, which is the time required to enter the vertiport after receiving the signal.
@@
A more profound understanding of these events is gleaned from the animation presented in Section 5.2. The root cause of most incidents lies in aircraft entering the space, particularly when other vehicles are near the gate. In an operational environment, this problem is solved by pre-arrival management, which informs aircraft already in the cruise phase about arrival times and provides advisories for in-time arrivals. Recognizing this observation, we conducted a secondary simulation study, incorporating a simple additional safety check for airspace entrance. Specifically, we examined whether any aircraft within the airspace maintained a Euclidean distance of less than $300\,\mathrm{m}$ to the respective gate. If affirmative, an alternative gate was selected for entrance; if all gates were occupied, no entrance is possible.
@@
In addition, Figures 7 and 8 show almost no false vertiport entrances except in cases with higher traffic density. Moreover, the airspace time increases linearly with the number of vehicles from approximately 220 s for $N=5$ to over 1000 s for $N=30$, which is primarily driven by the linear increase in the time to signal. This observation is expected since higher traffic density causes vehicles to wait longer for the entrance signal due to the limited capacity of the vertiport. Notably, the average entrance time decreases from approximately 45 s for $N=5$ vehicles to slightly above 30 s for $N=30$ vehicles. This reduction is due to the higher traffic density, which forces vehicles to spread out more broadly in the available airspace. Consequently, the distance of the closest vehicle to the eVTOL zone is reduced, leading to a quicker entrance time. However, this decrease in entrance time is negligible compared to the increase in the time to signal.
@@
6 Robustness analysis
In this section, we explore several methods to assess the robustness of Policy IV. First, we introduce Gaussian noise to the aircraft positions during decision-making to account for the limitations of practical sensing devices. Second, we modify the inbound air traffic to follow a Poisson distribution, reflecting common practices in conventional air traffic modeling. Third, we test Policy IV on small-scale drones to evaluate its transferability to real-world applications. Lastly, we include a discussion of the findings.
6.1 Positional noise
The location estimation of eVTOL vehicles is likely to rely on satellite-based radio navigation systems such as the Global Positioning System. After a vehicle determines its position, it can broadcast this information to surrounding aircraft, which use it as a basis for their distributed decision-making. However, practical sensing devices can introduce inaccuracies due to signal obstructions from surrounding objects, variations in atmospheric conditions, or limited receiver quality (Thin et al., 2016), possibly negatively impacting the overall operation.
We investigate how such positional noise affects the efficiency and safety of Policy IV. For vehicle $i$, we denote the ground-truth position vector at time $t$ as $p_{i,t}^{\text{G}}=\left(n_{i,t}^{\text{G}},e_{i,t}^{\text{G}}\right)^{\top}$, where $n_{i,t}^{\text{G}}$ and $e_{i,t}^{\text{G}}$ are the north and east components, respectively. Based on this, we construct the noisy position vector $p_{i,t}^{\text{N}}=\left(n_{i,t}^{\text{N}},e_{i,t}^{\text{N}}\right)^{\top}$ with north component $n_{i,t}^{\text{N}}$ and east component $e_{i,t}^{\text{N}}$ as follows:
$p_{i,t}^{\text{N}}=p_{i,t}^{\text{G}}+\epsilon_{i,t},$ (13)
where $\epsilon_{i,t}\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}\sigma_{\text{N}}^{2}&0\\ 0&\sigma_{\text{N}}^{2}\end{pmatrix}\right)$ is a realization of a two-dimensional zero-mean Gaussian random vector, with $\sigma_{\text{N}}$ being the standard deviation. The noise is assumed to be uncorrelated over time. In the following, the computations of vehicle observations (see Section 4.2), including distances, angles, and risk measures, rely entirely on the noisy positions $p_{i,t}^{\text{N}}$.
In particular, we consider three distinct noise intensities: Small noise with $\sigma_{\text{N}}=10\,\mathrm{m}$, medium noise with $\sigma_{\text{N}}=20\,\mathrm{m}$, and large noise with $\sigma_{\text{N}}=100\,\mathrm{m}$. We replicate the simulation study of Section 5.4 for each noise level to conduct a quantitative evaluation. The results are displayed in A. In addition, we provide an animation [^2] https://youtu.be/nVCsP6YOm58 to visualize the effect of the positional noise. In the animation, the ground truth positions are represented in darker colors, while the noisy positions used for action computation oscillate in lighter colors around the ground truth positions.
The animation and simulation results clearly demonstrate Policy IV’s robustness against positional measurement errors. The flight behavior remains smooth, maintaining high safety and efficiency levels for small to medium noise levels. We primarily attribute this strong performance to the architecture of the neural network in Figure 2. The temporal recurrent component effectively manages partial observability, ensuring Policy IV’s noise resilience. The incident frequency only increases in the large noise case, where accidents are also observed. However, we emphasize that $\sigma_{\rm N}=100\,\mathrm{m}$ constitutes an extreme scenario with limited practical relevance to truly test the limits of Policy IV. To summarize, our DRL-based approach offers strong robustness against measurement errors from real-world sensing devices.
6.2 Distribution of inbound traffic
The validation scenario of Section 5.3 considered several waves of aircraft, in which each wave consisted of four entering aircraft, one per gate. Furthermore, the simulation studies of Sections 5.4 and 6.1 assumed single aircraft entering at a random gate with a $15\,\mathrm{s}$ time gap to achieve a steady increase in traffic density. While these approaches are suitable for investigating the performance of our DRL policy, we acknowledge that inbound traffic in conventional air traffic is frequently modeled via a Poisson distribution (Lancia and Lulli, 2020). Therefore, we investigate Policy IV’s robustness against the distribution of inbound traffic via a scenario with Poisson-distributed arrivals. In particular, we assume all aircraft to enter the airspace via the south gate, representing a clustered arrival through one flight corridor. We assume four clusters with a $120\,\mathrm{s}$ time gap between consecutive clusters. Each cluster’s number of aircraft is sampled from a Poisson distribution with parameter $\lambda=5$ so that, in expectation, 20 aircraft are generated. The aircraft inside one cluster enter the airspace with a small time gap of $10\,\mathrm{s}$. Two independent runs of the scenario are shown in an animation [^3] https://youtu.be/FxuDZTx2kWw.
As the animation illustrates, the distribution of inbound traffic has minimal impact on operational performance, provided that the vehicles’ entrance does not immediately cause an incident, which can be mitigated with a simple safety check as discussed in Section 5.4. In the animation, the arriving Poisson sequence of vehicles quickly disperses in the airspace as they begin to spread out. Consequently, Policy IV demonstrates strong robustness to variations in the distribution of inbound traffic.
@@ 6.3 Sim-2-Real transfer
We demonstrate the real-world applicability of our DRL-based control approach by deploying Policy IV on five Crazyflie drones (Giernacki et al., 2017). The results are showcased in a video [^4] https://youtu.be/wbHUxARpHzM. As depicted in Figure 10, the Crazyflie is a nano quadcopter used as an experimental platform for research and education. One drone weighs approximately 27 g and offers battery for up to 7 min of flight time before recharging. The technical setup is as follows: The drones’ positions are estimated using a camera-based motion capture system, while the DRL-related computations are performed on a ThinkPad-P15v-Gen-3 notebook with an Intel i7-12700H processor. The ROS2-based Crazyswarm2 (IMRCLab, 2024) testbed acts as middleware for communicating with the drones. In particular, the notebook receives information from the motion capture system and computes the high-level control commands for each drone using the DRL policy. The latter provides heading commands, from which we compute a desired position for each drone assuming a constant velocity; see Section 3.1. These desired positions are sent to the drones, which are equipped with an STM32F405 onboard microcontroller to derive low-level actuator commands. More precisely, we use a differential flatness-based low-level controller (Mellinger and Kumar, 2011) to govern the drone actuators.
In our experimental setup, five drones are initially equidistantly spaced in the airspace. After a drone enters the vertiport, it is out of the scope of our DRL policy, and we hard-code the respective landing position to ensure there are no collisions. The radius of the airspace is set to $1.2\,\mathrm{m}$, and the drones maintain a velocity of approximately $0.15\,\mathrm{m/s}$. The video illustrates a stable and efficient operation, with the drones maintaining a safe distance and entering the vertiport sequentially. Crucially, we did not retrain the policy to accommodate the reduced size of the environment compared to the training scenarios. Since the agents’ observation space is properly normalized (see Section 4.2), adjusting the normalizing constants to match the smaller scale of the experiment suffices. The successful zero-shot adaptation of our policy is remarkable, considering that Sim-2-Real transfer often poses a significant challenge for RL algorithms (Zhao et al., 2020; Muratore et al., 2019).
@@
6.4 Discussion
The prior experiments and robustness analyses underscore the generalizability and practical usability of our DRL-based approach. In particular, we showcase that Policy IV is robust to partial observability in the form of noise and real-world disturbances. From a policy maker’s perspective, these observations make a strong case for DRL in contrast to traditional heuristic policies or local optimization routines. DRL excels in adapting to complex and dynamic environments where conventional methods may struggle due to their static nature. It can learn to handle uncertainty and stochasticity inherent in real-world environments more effectively (Kober et al., 2013; Mnih et al., 2015). Furthermore, DRL can consider multiple objectives simultaneously, in our case, minimizing collisions, guaranteeing vertiport entrance, staying in the airspace, and selecting comfortable actions. These trade-offs are often difficult to explicitly encode in a heuristic policy without sacrificing simplicity or performance. Lastly, from a computational point-of-view, our DRL approach has an extremely low inference time, allowing for real-time distributed decision-making with an arbitrary number of vehicles. To summarize, while heuristic policies have their place, particularly in well-defined and stable environments, DRL provides a promising approach for addressing the challenges and uncertainties inherent in future UAM systems.